OpenStudy (anonymous):

What is (m^3/4 times n^1/4)^2 over (m^1/4 times n^-1/2)^3 I tried to solve this problem but I am having a lot of trouble. Please explain as well so I can get a better understanding. Thank you so much.

OpenStudy (anonymous):

\[(m^{3/4} \times n^{1/4})^2 \over (m^{1/4} \times n^{-1/2})^3\] \[=m^{(3 \times2)/4} \times n^{(1\times 2)/4} \over m^{(1\times 3)/4} \times n^{(-1\times 3)/2}\] \[=m^{3/2} \times n^{1/2} \over m^{3/4} \times n^{-3/2} \] \[=m^{3/2-3/4} \times n^{1/2-(-3/2)} =m^{(6-3)/4} \times n^{1/2+3/2}=m^{3/4} \times n^{4/2}\] \[=m^{3/4} \times n^{2}\]

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

You are most welcome.